# footnotetext: * Equal contribution

### v1

@@ 3.2 Convergent Functions: Multi-Design Control Policies
We adopt the Transformer-based policy architecture (5 layers/2 heads) from [63] for our multi-design policies. Each policy is trained for 100 million simulation steps using a batch size of 5120 and an initial learning rate of $3\times 10^{-4}$ with cosine annealing. Training is distributed across six A40 GPUs, with each GPU handling 6–7 cluster-specific policies in parallel (more details in the Appendix).
@@ 3.3 Divergent Forms: Co-evolving Elites using Dynamic Local Search within Clusters
To overcome these limitations, we sample divergent forms using Dynamic Local Search (DLS) [25, 26, 27], combined with localized competition within each cluster. This allows clusters to independently evolve high-performing and behaviorally diverse morphologies (see different locomotion behaviors in Fig. 2). DLS promotes broader and more adaptive exploration within each cluster, reducing the risk of getting stuck in local minima. Our multi-design policy $\pi_{\theta}^{C_{k}}$ acts as a dynamic evaluation function, continually updated via PPO [32] on a dynamic pool of elite designs $\mathcal{P}_{i}^{C_{k}}$. This enables efficient assessment of diverse candidates sampled broadly across the cluster, rather than relying on narrowly mutated variants. Because evaluations are made efficient through policy sharing, we can use random search instead of mutation—allowing us to quickly filter out poor designs and retain promising ones.
@@
How to choose the number of clusters? Number of clusters plays a key role in balancing intra-cluster behavioral similarity with the breadth of localized competition. Within each cluster, morphologies should be behaviorally similar enough for the policy to generalize effectively, while the clustering should remain fine-grained enough to capture morphological diversity. With $N_{c}=10$, clusters roughly align with limb count, resulting in overly coarse groupings. We select $N_{c}=40$ to better capture finer morphological distinctions while maintaining a feasible computational budget. Each policy evaluates up to $\left\lfloor\frac{N_{\text{iter}}}{f_{\text{diff}}}\right\rfloor\cdot N_{\text{sample}}+N_{w}\sim 78{,}000$ samples during training, allowing each morphology to be evaluated approximately 8 times on average. These repeated evaluations are valuable, as early policies may be unreliable; re-evaluation with improved policies can uncover better-performing behaviors (see Appendix for details).
@@
Figure 5: Loki evolves significantly more sparse solutions (measured by the average distance to $k$-nearest neighbors). This is due to using Dynamic Local Search for sampling designs, rather than mutations. Higher sparseness indicates more diversity. (Statistical significance was assessed using independent samples t-tests; *$\text{*}P<0.005$; **$\text{**}P<10^{-11}*$; ***$\text{***}P<10^{-16}**$).
@@ 4 Experiments
Baselines and Experiment Details. We compare our method with DERL [4], a prior evolution-based co-design approach, and CVT-Map-Elites [22], a representative Quality-Diversity baseline. For a fair comparison, we evaluate the top $N=100$ final evolved morphologies (elites) from each algorithm. For methods using $N_{c}=40$ clusters, we select the top 2–3 agents from the final training pool of each cluster to form the final population. We include the following baselines and ablations:
@@
    DERL [4]: We use the publicly released set of $N=100$ agents evolved on flat terrain. DERL introduces a scalable framework for evolving morphologies in the UNIMAL design space using a bi-level optimization strategy: an outer evolutionary loop mutates designs, while an inner reinforcement learning loop trains a separate MLP policy for each agent for over 5 million steps. DERL employs distributed, asynchronous evolutionary search to parallelize this process, evolving and training 4,000 morphologies over an average of 10 generations to select the top 100.
@@
    MAP-Elites [23, 49]: A Quality-Diversity (QD) baseline where the morphology space is discretized into $N_{c}=40$ niches using K-means clustering on raw morphology parameters. Each cluster serves as a cell in the MAP-Elites archive maintaining 3 elites throughout evolution. The algorithm explores the same number of morphologies as DERL (40 generations $\times$ 100 morphologies per generation=4000), training a separate MLP policy for each agent for 5M steps.
@@ 4.1 Performance and Diversity of the Evolved Designs
The design space includes both fast and slow learning morphologies, requiring varying amounts of interaction to master locomotion on flat terrain. Training single-design controllers from scratch is constrained by limited budgets—e.g. DERL [4] trains each morphology using MLP policies for 5 million steps, which is often insufficient to learn more complex behaviors, such as spinning (see Appendix). Global competition causes the algorithm to favor faster learners like cheetah-like morphologies, reducing the diversity of behaviors. In contrast, Loki promotes both fast and slow learners by localizing competition within clusters, using powerful cluster-specific policies, and incorporating Dynamic Local Search (DLS) to encourage broader exploration within each cluster. This leads to greater behavioral and morphological diversity (Fig. 2).
In Tab. 2, we compare Loki for locomotion on flat terrain (FT) against baselines using three core metrics commonly used to evaluate Quality-Diversity (QD) algorithms [64, 19]: Maximum Fitness (the highest reward achieved by any agent in the final population), Coverage (the number of clusters filled by the final population), and QD-Score (sum of the highest fitness values discovered in each cluster, capturing both quality and diversity). For fair comparison of QD-score and coverage, we assign top 100 agents from each baseline to $N_{c}{=}40$ latent clusters (details in Appendix).
@@
Fig. 5 compares the sparseness  [20] of solutions as a novelty metric. A simple measure of sparseness $\rho$ at a point $x$ is $\rho(x)=\frac{1}{k}\sum_{i=1}^{k}\operatorname{dist}\left(x,\mu_{i}\right)$, where $\mu_{i}$ is the $i$th-nearest neighbor of $x$ in the morphology latent space. Loki achieves the highest sparseness among the three learning-based methods, closely approaching Random. This is largely due to Loki’s use of Dynamic Local Search (DLS) to sample diverse designs within each cluster, as opposed to the incremental mutations used in DERL and Latent-MAP-Elites, which tend to produce solutions in local neighborhoods.
@@ 4.2 Transitioning to New Tasks and Environments
Morphology-Level. Loki produces a diverse set of solutions that are better suited for adaptation to various unseen tasks and environments, reducing the need to re-evolve agents for each setting. We evaluate generalization across eight test tasks, each requiring a distinct set of skills. For each method, the final set of $N=100$ evolved morphologies (elites) is independently trained from scratch on each test task using MLP-based policies, with 5 random seeds and training durations of 5, 15, or 20 million steps depending on task difficulty. Obstacle (n=150) is more challenging with three times more obstacles than obstacle (n=50). In the bump task, agents must learn behaviors like jumping or climbing without explicit information about the bump’s location. All tasks except obstacle (n=150) and bump were originally introduced in [4] (more details in Appendix).
@@ 6 Limitations
  • Detailed description of the different test tasks (App.C)
  • Detailed description of the QD-score metric (App.D)
  • Description of how efficiency metrics in Tab.LABEL:tab:efficiency are calculated (App.E)
  • Pseudo-code for the Map-Elites baseline (App.F)
  • Full hyperparameter details for the experiments (App. G)
@@ Appendix B Learning Speed of the Evolved Morphologies (Fast & Slow Learners)
We examine the correlation between learning speed and the final performance of evolved agents. To evaluate performance across different training durations, we train single-agent MLP policies [4] for both 5M and 15M steps on flat terrain. This evaluation excludes agents from clusters with consistently low training rewards or predominantly simple morphologies. As shown in Fig. 9, the top-performing agents under short- and long-term training are largely disjoint, revealing the presence of both fast- and slow-learning morphologies. Notably, high short-term performance does not necessarily indicate high long-term performance—some slow learners achieve superior results after 15M steps despite underperforming at 5M steps.
@@
Appendix C Test Task Descriptions
We describe the task specifications and required skills for the eight test tasks used in our morphology-level evaluation. All tasks, except for Bump and Obstacle (n=150), were introduced in prior work [4]. We adopt the same task configurations and reward structures as outlined therein.
[…]
@@ Appendix C Test Task Descriptions
[…]
@@
Appendix D QD-Score (a quality-diversity metric)
QD-score [64, 19, 55] is a more comprehensive metric than maximum fitness, as it captures not only the performance of the single best agent but also the diversity of high-performing solutions across the search space. Given the vast combinatorial complexity of the UNIMAL space, QD-score is particularly well-suited for evaluating the quality and spread of evolved morphologies.
@@ Appendix D QD-Score (a quality-diversity metric)
[…]
@@
Appendix E Efficiency Metrics in Table LABEL:tab:efficiency.
@@
[…]
@@
[…]
@@
[…]
@@
Note that each morphology is evaluated approximately 8 times on average. These repeated evaluations are valuable because early-stage policies may be unreliable. Re-evaluating morphologies with progressively improved policies allows for the discovery of higher-performing behaviors.
FLOPs per searched morphology. Given the MLP and Transformer policy model architectures used in prior work [4, 63], we compute the training FLOPs per searched morphology as shown in Eq. E with the values in Tab. 4.
[…]
@@
Appendix F MAP-Elites
We provide the algorithm for the Map-Elites [23, 49] baseline in Alg. 2.
@@ Appendix F MAP-Elites
Appendix G Implementation Details
@@
Learning rate scheduler ReduceLROnPlateau [65]
@@
$[\beta_{\text{min}},\beta_{\text{max}}]$ $[10^{-5},10^{-2}]$

### v2

@@ 3.2 Convergent Functions: Multi-Design Control Policies
We adopt the Transformer-based policy architecture (5 layers/2 heads) from gupta2022metamorph for our multi-design policies. Each policy is trained for 100 million simulation steps using a batch size of 5120 and an initial learning rate of $3\times 10^{-4}$ with cosine annealing. Training is distributed across six A40 GPUs, with each GPU handling 6–7 cluster-specific policies in parallel (more details in the Appendix H).
@@ 3.3 Divergent Forms: Co-evolving Elites using Dynamic Local Search within Clusters
To overcome these limitations, we sample divergent forms using Dynamic Local Search (DLS) Hoos2015; Hutter2002ScalingAP; Pullan_2006, combined with localized competition within each cluster. This allows clusters to independently evolve high-performing and behaviorally diverse morphologies (see different locomotion behaviors in Fig. 2). DLS promotes broader and more adaptive exploration within each cluster, reducing the risk of getting stuck in local minima. Our multi-design policy $\pi_{\theta}^{C_{k}}$ acts as a dynamic evaluation function, continually updated via PPO ppo on a dynamic pool of elite designs $\mathcal{P}_{i}^{C_{k}}$. This enables efficient assessment of diverse candidates sampled broadly across the cluster, rather than relying on narrowly mutated variants. Because evaluations are made efficient through policy sharing, we can use random search instead of mutation—allowing us to quickly filter out poor designs and retain promising ones.
@@
How to choose the number of clusters? Number of clusters plays a key role in balancing intra-cluster behavioral similarity with the breadth of localized competition. Within each cluster, morphologies should be behaviorally similar enough for the policy to generalize effectively, while the clustering should remain fine-grained enough to capture morphological diversity. With $N_{c}=10$, clusters roughly align with limb count, resulting in overly coarse groupings. We select $N_{c}=40$ to better capture finer morphological distinctions while maintaining a feasible computational budget. Each policy evaluates up to $\left\lfloor\frac{N_{\text{iter}}}{f_{\text{diff}}}\right\rfloor\cdot N_{\text{sample}}+N_{w}\sim 78{,}000$ samples during training, allowing each morphology to be evaluated approximately 6 times on average. These repeated evaluations are valuable, as early policies may be unreliable; re-evaluation with improved policies can uncover better-performing behaviors (see Appendix for details).
@@
Figure 5: Loki evolves significantly more sparse solutions (measured by the average distance to $k$-nearest neighbors). This is due to using Dynamic Local Search for sampling designs, rather than mutations. Higher sparseness indicates more diversity. (Statistical significance was assessed using independent samples t-tests; $\text{*}P<0.005$; $\text{**}P<10^{-11}$; $\text{***}P<10^{-16}$).
@@ 4 Experiments
Baselines and Experiment Details. We compare our method with DERL gupta2021embodied, a prior evolution-based co-design approach, and CVT-Map-Elites Vassiliades2016UsingCV, a representative Quality-Diversity baseline. For a fair comparison, we evaluate the top $N=100$ final evolved morphologies (elites) from each algorithm. For methods using $N_{c}=40$ clusters, we select the top 2–3 agents from the final training pool of each cluster to form the final population. We include the following baselines and ablations:
@@
    DERL gupta2021embodied: We use the publicly released set of $N=100$ agents evolved on flat terrain. DERL introduces a scalable framework for evolving morphologies in the UNIMAL design space using a bi-level optimization strategy: an outer evolutionary loop mutates designs, while an inner reinforcement learning loop trains a separate MLP policy for each agent for over 5 million steps. DERL employs distributed, asynchronous evolutionary search to parallelize this process, evolving and training 4,000 morphologies over an average of 10 generations to select the top 100.
@@
    MAP-Elites Mouret2015IlluminatingSS; vassiliades2017using: A Quality-Diversity (QD) baseline where the morphology space is discretized into $N_{c}=40$ niches using K-means clustering on raw morphology parameters. Each cluster serves as a cell in the MAP-Elites archive maintaining 3 elites throughout evolution. The algorithm explores the same number of morphologies as DERL (40 generations $\times$ 100 morphologies per generation=4000), training a separate MLP policy for each agent for 5M steps.
@@ 4.1 Performance and Diversity of the Evolved Designs
The design space includes both fast and slow learning morphologies, requiring varying amounts of interaction to master locomotion on flat terrain. Training single-design controllers from scratch is constrained by limited budgets—e.g. DERL gupta2021embodied trains each morphology using MLP policies for 5 million steps, which is often insufficient to learn more complex behaviors, such as spinning (see Appendix B). Global competition causes the algorithm to favor faster learners like cheetah-like morphologies, reducing the diversity of behaviors. In contrast, Loki promotes both fast and slow learners by localizing competition within clusters, using powerful cluster-specific policies, and incorporating Dynamic Local Search (DLS) to encourage broader exploration within each cluster. This leads to greater behavioral and morphological diversity (Fig. 2).
In Tab. 2, we compare Loki for locomotion on flat terrain (FT) against baselines using three core metrics commonly used to evaluate Quality-Diversity (QD) algorithms flageat2022benchmarking; pugh2016quality: Maximum Fitness (the highest reward achieved by any agent in the final population), Coverage (the number of clusters filled by the final population), and QD-Score (sum of the highest fitness values discovered in each cluster, capturing both quality and diversity). For fair comparison of QD-score and coverage, we assign top 100 agents from each baseline to $N_{c}{=}40$ latent clusters (details in Appendix E).
@@
Fig. 5 compares the sparseness  Lehman2011EvolvingAD of solutions as a novelty metric. A simple measure of sparseness $\rho$ at a point $x$ is $\rho(x)=\frac{1}{k}\sum_{i=1}^{k}\operatorname{dist}\left(x,\mu_{i}\right)$, where $\mu_{i}$ is the $i$th-nearest neighbor of $x$ in the morphology latent space. Loki achieves the highest sparseness among the three learning-based methods, closely approaching Random. This is largely due to Loki’s use of Dynamic Local Search (DLS) to sample diverse designs within each cluster, as opposed to the incremental mutations used in DERL and Latent-MAP-Elites, which tend to produce solutions in local neighborhoods.
@@ 4.2 Transitioning to New Tasks and Environments
Morphology-Level. Loki produces a diverse set of solutions that are better suited for adaptation to various unseen tasks and environments, reducing the need to re-evolve agents for each setting. We evaluate generalization across eight test tasks, each requiring a distinct set of skills. For each method, the final set of $N=100$ evolved morphologies (elites) is independently trained from scratch on each test task using MLP-based policies, with 5 random seeds and training durations of 5, 15, or 20 million steps depending on task difficulty. Obstacle (n=150) is more challenging with three times more obstacles than obstacle (n=50). In the bump task, agents must learn behaviors like jumping or climbing without explicit information about the bump’s location. All tasks except obstacle (n=150) and bump were originally introduced in gupta2021embodied (more details in Appendix D).
@@ 6 Limitations
  • Procedure for sampling random morphologies within the UNIMAL space (App.C)
  • Detailed description of the different test tasks (App.D)
  • Detailed description of the QD-score metric (App.E)
  • Description of how efficiency metrics in Tab.LABEL:tab:efficiency are calculated (App.F)
  • Pseudo-code for the Map-Elites baseline (App.G)
  • Full hyperparameter details for the experiments (App. H)
  • Broader Societal Impacts (App. I)
@@ Appendix B Learning Speed of the Evolved Morphologies (Fast & Slow Learners)
We examine the correlation between learning speed and the final performance of evolved agents. To evaluate performance across different training durations, we train single-agent MLP policies [gupta2021embodied] for both 5M and 15M steps on flat terrain. This evaluation excludes agents from clusters with consistently low training rewards or predominantly simple morphologies. As shown in Fig. 9, the top-performing agents under short- and long-term training are largely disjoint, revealing the presence of both fast- and slow-learning morphologies. Notably, high short-term performance does not necessarily indicate high long-term performance—some slow learners achieve superior results after 15M steps despite underperforming at 5M steps.
@@
Appendix C Sampling Random Morphologies within the UNIMAL Space
The process of sampling random morphologies follows the procedure introduced in prior work [gupta2021embodied]. During the population initialization phase, a new morphology is created by first sampling the total number of limbs to grow, followed by a series of mutation operations until the desired number of limbs is reached. These mutations include: growing or deleting limbs, mutating limb parameters, density, degrees of freedom (DoF), gear ratios, and joint angles. For each mutation, the parameters are uniformly sampled from predefined ranges specified in [gupta2021embodied]. The key difference is that DERL [gupta2021embodied] samples parameter values from discrete sets, while LOKI samples continuously within each range, enabling a denser and more comprehensive coverage of the morphology space. Table 4 presents the parameter ranges used to create our morphologies. The notation range(a, b) denotes a continuous range from a to b.
Parameters Sampling Range
Max limbs 11
Limb radius range(0.02, 0.06)
Limb height range(0.2, 0.4)
Limb density range(500, 1000)
Limb orientation $\theta$ [0, 45, 90, 135, 180, 225, 270, 315]
Limb orientation $\phi$ [90, 135, 180]
Head radius 0.10
Head density range(500, 1000)
Joint axis [x, y, xy]
Motor gear range range(150, 300)
Joint limits [(-30, 0), (0, 30), (-30, 30), (-45, 45), (-45, 0), (0, 45),
(0, -60), (0, 60), (-60, 60), (-90, 0), (0, 90), (-60, 30), (-30, 60)]
Table 4: Design parameters used for sampling random morphologies in the UNIMAL space.
Appendix D Test Task Descriptions
We describe the task specifications and required skills for the eight test tasks used in our morphology-level evaluation. All tasks, except for Bump and Obstacle (n=150), were introduced in prior work [gupta2021embodied]. We adopt the same task configurations and reward structures as outlined therein.
Patrol. The agent must repeatedly traverse between two target points separated by $10\text{m}$ along the x-axis. High performance in this task requires rapid acceleration, short bursts of speed, and quick directional changes. (Training steps: 5 million)
@@ Appendix D Test Task Descriptions
Push Box Incline. The agent must push a box with a side length of $0.2\text{m}$ up an inclined plane. The environment is a $80\times 40\text{m}^{2}$ rectangular arena tilted at a 10-degree angle. The agent starts at one end of the arena and is tasked with propelling the box forward along the slope. (Training steps: 5 million)
Obstacle (n=50, 150). The agent must navigate through a cluttered environment filled with static obstacles to reach the end of the arena. Each box-shaped obstacle has a width and length between $0.5\text{m}$ and $3~\text{m}$, with a fixed height of $2~\text{m}$. n denotes the number of randomly distributed obstacles across a flat $150\times 60\text{m}^{2}$ terrain. (Training steps: 5 million)
Bump. The agent must traverse an arena filled with 250 low-profile obstacles randomly placed on a flat $150\times 60\text{m}^{2}$ terrain. Each obstacle has a width and length between $0.8\text{m}$ and $1.6\text{m}$, and a height between $0.1\text{m}$ and $0.25\text{m}$. As obstacle height is comparable to the agent’s body, this task promotes behaviors such as jumping or climbing, adding complexity to locomotion and body coordination. (Training steps: 15 million)
Manipulate Ball. The agent must move a ball from a random source location to a fixed target. A ball of radius $0.2~\text{m}$ is placed at a random location in a flat, square $30\times 30\text{m}^{2}$ arena, with the agent initialized at the center. This task requires a fine interplay of locomotion and object manipulation, as the agent must influence the ball’s motion through contact while maintaining its own balance and stability. (Training steps: 20 million)
@@
Appendix E QD-Score (a quality-diversity metric)
QD-score [flageat2022benchmarking, pugh2016quality, nordmoen2021map] is a more comprehensive metric than maximum fitness, as it captures not only the performance of the single best agent but also the diversity of high-performing solutions across the search space. Given the vast combinatorial complexity of the UNIMAL space, QD-score is particularly well-suited for evaluating the quality and spread of evolved morphologies.
@@ Appendix E QD-Score (a quality-diversity metric)
$\text{QD-score}=\frac{1}{N_{c}}\sum_{i=1}^{N_{c}}\mathds{1}_{\|\mathcal{M}_{i}\|>0}\cdot\frac{f_{i}-f_{\text{min}}}{f_{\text{max}}-f_{\text{min}}}$ (1)
@@
Appendix F Efficiency Metrics in Table LABEL:tab:efficiency.
@@
Total interactions $=N_{c}\cdot\left(\text{\# interactions {\scriptsize(per cluster)}}+\text{\# evaluated samples}\cdot\text{episode length}\right)$
@@
In contrast, DERL trains 4,000 agents independently using a single-agent MLP policy, with each agent trained for $5\text{M}$ environment steps—resulting in a total of $20\text{B}$ interactions. Since DERL uses the training reward directly for agent selection, no separate evaluation phase is required.
@@
# $\text{\# of covered morphologies {\scriptsize(per cluster)}}\leq\left\lfloor\frac{N_{\text{iter}}}{f_{\text{diff}}}\right\rfloor\cdot N_{\text{sample}}+N_{w}\approx 78{,}000$ (3)
# of searched morphologies =# $=\text{\# of covered morphologies {\scriptsize(per cluster)}}\cdot N_{c}$
@@
Note that each morphology is evaluated approximately 6 times on average. These repeated evaluations are valuable because early-stage policies may be unreliable. Re-evaluating morphologies with progressively improved policies allows for the discovery of higher-performing behaviors.
FLOPs per searched morphology. Given the MLP and Transformer policy model architectures used in prior work [gupta2021embodied, gupta2022metamorph], we compute the training FLOPs per searched morphology as shown in Eq. F with the values in Tab. 5.
FLOPs (per step) $=2\times\text{FLOPs {\scriptsize(per forward pass)}}\times\text{Batch size}$
FLOPs (per model) $=\text{FLOPs {\scriptsize(per step)}}\times\text{PPO epochs}\times\text{(\# of iterations)}$
FLOPs (per morphology) $=\text{FLOPs {\scriptsize(per model)}}/\text{(\# of searched morphologies)}$ (5)
@@
Appendix G MAP-Elites
We provide the algorithm for the Map-Elites [Mouret2015IlluminatingSS, vassiliades2017using] baseline in Alg. 2.
@@ Appendix G MAP-Elites
Appendix H Implementation Details
@@
Learning rate scheduler ReduceLROnPlateau [paszke2019pytorchimperativestylehighperformance]
@@
$\beta_{\text{min}},\beta_{\text{max}}] [$$10^{-5},10^{-2}]$
@@
Appendix I Societal Impacts
Positive impacts: Our work aims to challenge prevailing assumptions about optimal robot morphology, potentially reshaping how robotic design is approached. By promoting diversity and functionality beyond conventional forms, LOKI encourages exploration of unconventional yet effective morphologies. We believe this contributes toward a more inclusive and biologically inspired understanding of embodiment, potentially serving as a bridge between AI, robotics, and the natural sciences.
Negative impacts: A potential concern is that the ability to autonomously generate a large number of novel and capable morphologies could be misused in contexts that prioritize performance over safety. For instance, this approach could enable rapid prototyping of morphologies for autonomous systems without adequate human oversight, increasing the risk of deploying untested designs in sensitive environments.